OpenStudy (anonymous):

Three vertices of a rectangle are given. Find the coordinates of the fourth vertex. (2, 5), (6, 1), (4, –1)

OpenStudy (anonymous):

i will give medals pleez help

OpenStudy (anonymous):

i really need help i will become a fan

OpenStudy (er.mohd.amir):

mid point of AC is (3,2) now mid point BD is if D is(h,k) is (6+h/2,1+k/2) which is equal mid point of AC (6+h)/2=3 6+h=6 h=0 1+k)/2=2 1+k=4 k=3 point (0,3)